f(x)5x-7 and g(x)=5-x^2
a f(g(-1))
b g(f(-1)

Answer by jim_thompson5910(35256)

I'll do part a) to get you started

f(x) = 5x-7

f(g(x)) = 5(g(x))-7 ... replace every x with g(x)

f(g(x)) = 5(5-x^2)-7 ... replace g(x) on the right side with 5-x^2 (valid because g(x) = 5-x^2)

f(g(x)) = 5(5-x^2)-7

f(g(x)) = 25-5x^2-7

f(g(x)) = -5x^2+18

Now plug in x = -1

f(g(x)) = -5x^2+18

f(g(-1)) = -5(-1)^2+18 ... replace every x with -1

f(g(-1)) = -5(1)+18

f(g(-1)) = -5+18

f(g(-1)) = 13

I leave part b) for you.